# EEGDM: Learning EEG Representation with Latent Diffusion Model

### v1

@@ Abstract
While electroencephalography (EEG) signal analysis using deep learning has shown great promise, existing approaches still face significant challenges in learning generalizable representations that perform well across diverse tasks, particularly when training data is limited. Current EEG representation learning methods including EEGPT and LaBraM typically rely on simple masked reconstruction objective, which may not fully capture the rich semantic information and complex patterns inherent in EEG signals. In this paper, we propose EEGDM, a novel self-supervised EEG representation learning method based on the latent diffusion model, which leverages EEG signal generation as a self-supervised objective, turning the diffusion model into a strong representation learner capable of capturing EEG semantics. EEGDM incorporates an EEG encoder that distills EEG signals and their channel augmentations into a compact representation, acting as conditional information to guide the diffusion model for generating EEG signals. This design endows EEGDM with a compact latent space, which not only offers ample control over the generative process but also can be leveraged for downstream tasks. Experimental results show that EEGDM (1) can reconstruct high-quality EEG signals, (2) effectively learns robust representations, and (3) achieves competitive performance with modest pre-training data size across diverse downstream tasks, underscoring its generalizability and practical utility.
Introduction
@@
Electroencephalography (EEG) is a non-invasive technique that measures spontaneous electrical activity in the brain through scalp electrodes. The interpretation and utilization of EEG signals generally require sophisticated signal processing pipelines and advanced feature extraction methodologies (Mohsenvand et al. 2020). Conventional EEG signal analysis typically employs manually engineered operations tailored to specific EEG features of interest (e.g., band-pass filtering in a specific frequency range), potentially excluding other informative signal components (Zander and Kothe 2011).
@@
Recently, deep learning has introduced a new research paradigm for EEG signal analysis due to its capability to automatically learn efficient representations from original data (Zhao, Yan, and Lu 2021; Zheng et al. 2023; Chen et al. 2025). Current deep learning-based approaches, which predominantly rely on large-scale EEG datasets, can be broadly classified into two categories: (1) task-specific models, and (2) multi-task large models. Task-specific models face significant challenges in generalization due to substantial variations across datasets in channel configurations, sampling rates, and EEG recording time lengths. These discrepancies necessitate model training on individual datasets, thereby severely limiting cross-dataset generalizability. For multiple tasks, two large EEG models, EEGPT (Wang et al. 2024) and LaBraM (Jiang, Zhao, and Lu 2024) have recently been proposed. Both models learn generic representations through masked reconstruction pre-training objective before being fine-tuned for various downstream tasks. However, LaBraM uses a substantially larger dataset, and EEGPT’s performance across diverse specific tasks still leaves room for improvement.
@@
In this paper, we introduce the Diffusion Model (DM) (Ho, Jain, and Abbeel 2020; Dhariwal and Nichol 2021) into EEG representation learning and propose a novel self-supervised framework called EEGDM. Our framework leverages EEG signal synthesis as a self-supervised objective, turning the diffusion model into a strong representation learner capable of capturing EEG semantics. Unlike popular masked reconstruction approaches that primarily focus on local feature recovery, DM learns effective representations through a self-supervised denoising objective that requires understanding the underlying data distribution. This makes DM potentially valuable for generic EEG representation learning. Building on this observation, our framework uses an EEG encoder that transforms input signals and their channel augmentations into compact latent representations, which condition the diffusion process. This design enables the encoder to learn the effective and generic representations inherent in EEG signals that are crucial for downstream tasks.
@@
We note that EEG signals present unique challenges due to their inherently low signal-to-noise ratio (SNR) and complex characteristics, including stochasticity, nonstationarity, and nonlinearity (Gramfort et al. 2013; Cole and Voytek 2019). To address these challenges, we incorporate two key technical components in our EEGDM framework. First, we employ channel augmentation that provides diverse training perspectives while enriching the conditional information. Second, we utilize Principal Component Analysis (PCA) to project EEG signals into a latent space with an improved SNR, as it enhances signal-relevant components while suppressing noise-related ones (Sanei and Chambers 2013). This transformation mitigates the challenge of reconstructing the original noisy signal. Inspired by the success of state-of-the-art DMs (e.g., Latent Diffusion Models (LDMs) (Rombach et al. 2022) and Diffusion Transformers (DiTs) (Peebles and Xie 2023)) operating in latent spaces, we apply our DM to this PCA-derived latent space rather than the original noisy EEG space. This enables EEGDM to focus on learning meaningful signal patterns for effective representation learning across diverse downstream tasks. Our experiments show that EEGDM effectively learns robust representations, achieving competitive performance across diverse downstream tasks with modest pre-training data requirements and exhibiting strong generalizability.
@@
  • We propose EEGDM, a novel diffusion model-based self-supervised framework for EEG representation learning. By leveraging EEG signal generation as the self-supervised objective, EEGDM transforms the diffusion model into a powerful representation learner that captures rich EEG semantics, even with limited pre-training data.
@@
  • We validate the effectiveness and generalizability of EEGDM across multiple EEG analysis tasks, demonstrating its (a) competitive performance compared to existing methods and (b) strong cross-dataset transferability and practical utility in EEG applications.
Related Work
Self-supervised EEG Representation Learning
Self-supervised learning can leverage large amounts of unlabeled data for pre-training, thereby reducing the reliance on manual annotations. This methodology has led to significant breakthroughs in natural language processing and computer vision (Kenton and Toutanova 2019; He et al. 2022). However, its full potential in EEG analysis remains largely unexplored (Jiang et al. 2025).
Some studies (Mohsenvand et al. 2020; Banville et al. 2021) have explored learning representations of EEG signals via contrastive learning, which trains a channel feature extractor by extending the SimCLR framework (Chen et al. 2020) to time series data. Another line of work extends the mask idea to EEG analysis (Mohammadi Foumani et al. 2024). For instance, BENDR (Kostas et al. 2021) uses a convolutional encoder to extract EEG features from local time windows, masks a portion of them, and utilizes a Transformer to predict the information in the masked part. LaBraM (Jiang, Zhao, and Lu 2024) pre-trains a Transformer model by predicting the original neural codes of masked EEG channel patches based on a defined neural codebook. EEGPT (Wang et al. 2024) introduces the alignment of spatio-temporal representation, implemented by a mask-based dual self-supervised learning for efficient extraction of EEG features.
@@
Diffusion Model
Diffusion models learn complex data distributions through a progressive denoising process that requires a deep understanding of the underlying data structure and semantics. This fundamental characteristic endows them with dual capabilities: generating high-quality samples and learning meaningful representations inherent in the data distribution. The pioneering work by Sohl-Dickstein et al. (Sohl-Dickstein et al. 2015) and subsequent advances by Ho et al. (Ho, Jain, and Abbeel 2020) established diffusion models as powerful tools, leading to widespread applications in generative tasks (Rombach et al. 2022; Peebles and Xie 2023; Zhang, Rao, and Agrawala 2023). More recently, their potential for representation learning has gained attention, with emerging studies showing promising results in learning discriminative features (Mittal et al. 2023; Chen et al. 2024b; Hudson et al. 2024).
In the EEG domain, existing diffusion-based approaches have primarily focused on generative applications, such as EEG synthesis and data augmentation (Huang et al. 2024; Chen et al. 2024a; Huang, Li, and Chen 2025), or a few specific supervised tasks (Kim et al. 2023, 2024). However, the potential of diffusion models for generic EEG representation learning remains largely unexplored. This presents a significant opportunity, as learning the complex temporal and spatial characteristics of EEG signals can benefit from the understanding of rich data distributions inherent in diffusion models’ denoising process. In this paper, we present a novel method that exploits the representational capacity of diffusion models for self-supervised EEG representation learning.
Method
@@
Framework Architecture
Our proposed self-supervised EEG representation learning framework, EEGDM, is formulated as a latent diffusion model (DM) conditioned on EEG channel augmentation. Our goal is to generate high-quality EEG signals while equipping the model with a generalizable understanding of EEG data that can be effectively adapted to various downstream tasks. It is worth noting that in the denoising process of DM, directly reconstructing raw EEG signals is challenging due to their inherently low SNR. To address this, we utilize PCA to preprocess the original EEG for noise reduction, and perform the reconstruction on the preprocessed EEG data with a relatively high SNR. The overall architecture of EEGDM is illustrated in Figure 1.
Specifically, given the original multi-channel EEG signals $X\in\mathbb{R}^{C\times T}$, where $C$ is the number of EEG electrodes (channels) and $T$ is the total timestamp, we segment $X$ into ${\lfloor{\frac{T-t^{s}}{s^{t}}}\rfloor}+1$ samples. This segmentation assumes a timestamp length [^1] Here superscript is used; subscript is reserved for diffusion formation in subsequent sections. $t^{s}$ per sample and a stride $s^{t}$. As shown in Figure 1, each multi-channel sample $x\in\mathbb{R}^{C\times{t^{s}}}$ is processed by PCA to obtain the latent representation $z$. Our diffusion model operates on this latent space, where noise is added and predicted. The denoised latent representation predicted by the diffusion model is projected back to the original space by inverse PCA to obtain the target EEG signal. Owing to the effectiveness of the DiT model in the latent space, we utilize it to predict the denoised latent representations. Furthermore, we introduce an EEG encoder that transforms the input source signal $x^{\prime}$ (including original EEG sample and its augmentations) into an EEG representation containing rich semantic information. The sum of this EEG representation and the noise timestep representation (i.e., timestep $t$ in the noise-adding process) encoded by the time embedding layer serves as the conditional input information for the diffusion model. This diffusion model design can utilize the semantically rich external EEG representation for the generative task, leading to a significant improvement in generative performance. Consequently, this enhances the model’s representation learning capability.
@@
The core motivation behind the above EEG encoder design arises from the inherently under-determined nature of the denoising task. The denoiser tends to exploit any relevant knowledge or informative cues to infer and recover the original data distribution as accurately as possible. This incentivizes the encoder to distill salient commonalities between the source and target EEG signals into a compact representation. We constrain the encoded EEG representation to a low-dimensional latent space and guide the denoiser through feature modulation. This strategy encourages the encoder to capture high-level semantic information in the EEG signals, while leaving the reconstruction of local details to the denoiser. Rather than reconstruction from scratch as most conventional denoisers do (Hinton and Salakhutdinov 2006; Higgins et al. 2017), our strategy enables the encoder to learn a representation that supports latent space denoising. This alleviates the burden of compressing all signal information into the representation, allowing it to prioritize the most distinctive and informative features of the EEG. Such a representation is particularly valuable for adapting to a wide range of downstream EEG tasks.
@@
Our EEG encoder is based on the Vision Transformer (ViT) architecture (Dosovitskiy 2021) that operates on the sequences of tokens. We segment each EEG channel into tokens and then flatten the tokens to form consistent biosignal sequences. We retain many of ViT’s best practices, enabling the EEG encoder to be a Transformer architecture that can process input EEG signals with various formats (different sampling rates, number of channels, and time lengths). The input of the encoder is the source EEG signal $x^{\prime}$, which consists of the original samples and the EEG channel augmentations. EEG augmentations are a set of transformations that preserve the semantic information in the EEG channels without altering the interpretation of the EEG data. Here, we employ two augmentations (zero-masking and amplitude scaling) which were shown in prior work (Mohsenvand et al. 2020) to be the most effective for extracting useful features.
@@
Specifically, we use a $\omega$-length window without overlap to segment each EEG channel sample into patches ($x^{\prime}=\{x^{\prime}_{j,k}\in\mathbb{R}^{\omega}|j=1,2,\cdots,C,k=1,2,\cdots,{\lfloor{\frac{t^{s}}{\omega}}\rfloor}\}$). The total number of patches for $x^{\prime}$ is ${\left|x^{\prime}\right|}=C{\lfloor{\frac{t^{s}}{\omega}}\rfloor}$. Then a temporal encoder (i.e., a temporal convolution block) is used to encode each EEG patch into a patch embedding $e^{\prime}=\{e^{\prime}_{j,k}\in\mathbb{R}^{d}|j=1,2,\cdots,C,k=1,2,\cdots,{\lfloor{\frac{t^{s}}{\omega}}\rfloor}\}$, where $d$ is the dimension of the embeddings. The learnable positional embeddings are added to all input patch embeddings. Finally, the sequence of embeddings is fed directly into the Transformer network. We use average pooling on the output embeddings to obtain the EEG representation, which is then combined with the noise timestep representation to guide the denoising of the diffusion model. In addition, we perform downstream EEG tasks by using task-specific prediction heads on the EEG representation. As shown in Figure 2, we train a linear classifier as the prediction head based on the pre-trained encoder in fine-tuning and then analyze the model’s performance on the downstream tasks, in which we employ cross-entropy loss. The implementation details can be found in the supplementary material.
EEG Data Generation by Diffusion Model

### v2

@@ Abstract
Recent advances in self-supervised learning for EEG representation have largely relied on masked reconstruction, where models are trained to recover randomly masked signal segments. While effective at modeling local dependencies, such objectives are inherently limited in capturing the global dynamics and long-range dependencies essential for characterizing neural activity. To address this limitation, we propose EEGDM, a novel self-supervised framework that leverages latent diffusion models to generate EEG signals as an objective. Unlike masked reconstruction, diffusion-based generation progressively denoises signals from noise to realism, compelling the model to capture holistic temporal patterns and cross-channel relationships. Specifically, EEGDM incorporates an EEG encoder that distills raw signals and their channel augmentations into a compact representation, acting as conditional information to guide the diffusion model for generating EEG signals. This design endows EEGDM with a compact latent space, which not only offers ample control over the generative process but also can be leveraged for downstream tasks. Experimental results show that EEGDM (1) reconstructs high-quality EEG signals, (2) learns robust representations, and (3) achieves competitive performance across diverse downstream tasks, thus exploring a new direction for self-supervised EEG representation learning.
Index Terms:
Self-supervised learning, diffusion model, EEG representation learning.
I Introduction
@@
Electroencephalography (EEG) is a non-invasive technique that measures spontaneous electrical activity in the brain through scalp electrodes. The interpretation and utilization of EEG signals generally require sophisticated signal processing pipelines and advanced feature extraction methodologies [mohsenvand2020contrastive]. Conventional EEG signal analysis typically employs manually engineered operations tailored to specific EEG features of interest (e.g., band-pass filtering in a specific frequency range), potentially excluding other informative signal components [zander2011towards].
@@
Recently, deep learning has introduced a new research paradigm for EEG signal analysis due to its capability to automatically learn efficient representations from original data [zhao2021plug, zheng2023scatterformer, chen2025long]. Current deep learning-based approaches, which predominantly rely on large-scale EEG datasets, can be broadly classified into two categories: (1) task-specific models, and (2) multi-task large models. Task-specific models face significant challenges in generalization due to substantial variations across datasets in channel configurations, sampling rates, and EEG recording time lengths. These discrepancies necessitate model training on individual datasets, thereby severely limiting cross-dataset generalizability. For multiple tasks, three large EEG models, EEGPT [wangeegpt], LaBraM [jiang2024large], and CBraMod [wang2024cbramod] have recently been proposed. These models learn generic representations through masked reconstruction pre-training objective before being fine-tuned for various downstream tasks. However, masked reconstruction has limited capacity in capturing the global dynamics and long-term dependencies essential for characterizing neural activity.
@@
In this paper, we introduce the Diffusion Model (DM) [ho2020denoising, dhariwal2021diffusion] into EEG representation learning and propose a novel self-supervised framework called EEGDM. Our framework leverages EEG signal synthesis as a self-supervised objective, turning the diffusion model into a strong representation learner capable of capturing EEG semantics. Unlike popular masked reconstruction approaches that primarily focus on local feature recovery, DM learns effective representations through a self-supervised denoising objective that requires understanding the underlying data distribution [ho2020denoising]. This makes DM potentially valuable for generic EEG representation learning. Building on this observation, our framework uses an EEG encoder that transforms input signals and their channel augmentations into compact latent representations, which condition the diffusion process. This design enables the encoder to learn the effective and generic representations inherent in EEG signals that are crucial for downstream tasks.
@@
We note that EEG signals present unique challenges due to their inherently low signal-to-noise ratio (SNR) and complex characteristics, including stochasticity, nonstationarity, and nonlinearity [gramfort2013time, cole2019cycle]. To address these challenges, we incorporate two key technical components in our EEGDM framework. First, we employ channel augmentation that provides diverse training perspectives while enriching the conditional information. Second, we utilize Principal Component Analysis (PCA) to project EEG signals into a latent space with an improved SNR, as it enhances signal-relevant components while suppressing noise-related ones [sanei2013eeg]. This transformation mitigates the challenge of reconstructing the original noisy signal. Inspired by the success of state-of-the-art DMs (e.g., Latent Diffusion Models (LDMs) [rombach2022high] and Diffusion Transformers (DiTs) [peebles2023scalable]) operating in latent spaces, we apply our DM to this PCA-derived latent space rather than the original noisy EEG space. This enables EEGDM to focus on learning meaningful signal patterns for effective representation learning across diverse downstream tasks. Our experiments show that EEGDM effectively learns robust representations, achieving competitive performance across diverse downstream tasks and exhibiting strong generalizability.
@@
  • We propose EEGDM, a novel diffusion model-based self-supervised framework for EEG representation learning. By leveraging EEG signal generation as the self-supervised objective, EEGDM transforms the diffusion model into a powerful representation learner that captures rich EEG semantics.
@@
  • We validate the effectiveness and generalizability of EEGDM across multiple EEG analysis tasks, demonstrating its (1) competitive performance compared to existing methods and (2) strong cross-dataset transferability and practical utility in EEG applications.
The remainder of this work is organized as follows. First, we describe the related works in Section II. Next, the proposed algorithm is introduced in detail in Section III. After that, experimental results are shown in Section IV. Finally, the conclusion is given in Section V.
II Related work
II-A Self-supervised EEG Representation Learning
Self-supervised learning can leverage large amounts of unlabeled data for pre-training, thereby reducing the reliance on manual annotations. This methodology has led to significant breakthroughs in natural language processing and computer vision [kenton2019bert, he2022masked]. However, its full potential in EEG analysis remains largely unexplored [jiang2024neurolm].
Some studies [mohsenvand2020contrastive, banville2021uncovering] have explored learning representations of EEG signals via contrastive learning, which trains a channel feature extractor by extending the SimCLR framework [chen2020simple] to time series data. Another line of work extends the mask idea to EEG analysis [mohammadi2024eeg2rep]. For instance, Chien et al. [chien2022maeeg] propose a masked auto-encoder, which learns the EEG representation by learning to reconstruct the masked EEG features using a Transformer architecture. BENDR [kostas2021bendr] uses a convolutional encoder to extract EEG features from local time windows, masks a portion of them, and utilizes a Transformer to predict the information in the masked part. LaBraM [jiang2024large] pre-trains a Transformer model by predicting the original neural codes of masked EEG channel patches based on a defined neural codebook. EEGPT [wangeegpt] introduces the alignment of spatio-temporal representation, implemented by a mask-based dual self-supervised learning for efficient extraction of EEG features. CBraMod [wang2024cbramod] introduces a structurally aware EEG foundation model that separately captures heterogeneous spatial–temporal dependencies through a criss-cross transformer and achieves strong cross-dataset generalization via large-scale masked pre-training.
@@ II-A Self-supervised EEG Representation Learning
II-B Diffusion Model
Diffusion models learn complex data distributions through a progressive denoising process that requires a deep understanding of the underlying data structure and semantics. This fundamental characteristic endows them with dual capabilities: generating high-quality samples and learning meaningful representations inherent in the data distribution. The pioneering work by Sohl-Dickstein et al. [sohl2015deep] and subsequent advances by Ho et al. [ho2020denoising] established diffusion models as powerful tools, leading to widespread applications in generative tasks [rombach2022high, peebles2023scalable, zhang2023adding]. More recently, their potential for representation learning has gained attention, with emerging studies showing promising results in learning discriminative features [mittal2023diffusion, chen2024deconstructing, hudson2024soda].
In the EEG domain, existing diffusion-based approaches have primarily focused on generative applications, such as EEG synthesis and data augmentation [huang2024eegdfus, chen2024improving, huang2025sad], or a few specific supervised tasks [kim2023diff, kim2024brain]. However, the potential of diffusion models for generic EEG representation learning remains largely unexplored. This presents a significant opportunity, as learning the complex temporal and spatial characteristics of EEG signals can benefit from the understanding of rich data distributions inherent in diffusion models’ denoising process. In this paper, we present a novel method that exploits the representational capacity of diffusion models for self-supervised EEG representation learning.
III Method
@@
III-A Framework Architecture
Our proposed self-supervised EEG representation learning framework, EEGDM, is formulated as a latent diffusion model (DM) conditioned on EEG channel augmentation. Our goal is to generate high-quality EEG signals while equipping the model with a generalizable understanding of EEG data that can be effectively adapted to various downstream tasks. It is worth noting that in the denoising process of DM, directly reconstructing raw EEG signals is challenging due to their inherently low SNR. To address this, we utilize PCA to preprocess the original EEG for noise reduction, and perform the reconstruction on the preprocessed EEG data with a relatively high SNR. The overall architecture of EEGDM is illustrated in Fig. 1.
Specifically, given the original multi-channel EEG signals $X\in\mathbb{R}^{C\times T}$, where $C$ is the number of EEG electrodes (channels) and $T$ is the total timestamp, we segment $X$ into ${\lfloor{\frac{T-t^{s}}{s^{t}}}\rfloor}+1$ samples. This segmentation assumes a timestamp length [^1] Here superscript is used; subscript is reserved for diffusion formation in subsequent sections. $t^{s}$ per sample and a stride $s^{t}$. As shown in Fig. 1, each multi-channel sample $x\in\mathbb{R}^{C\times{t^{s}}}$ is processed by PCA to obtain the latent representation $z$. Our diffusion model operates on this latent space, where noise is added and predicted. The denoised latent representation predicted by the diffusion model is projected back to the original space by inverse PCA to obtain the target EEG signal. Owing to the effectiveness of the DiT model in the latent space, we utilize it to predict the denoised latent representations. Furthermore, we introduce an EEG encoder that transforms the input source signal $x^{\prime}$ (including original EEG sample and its augmentations) into an EEG representation containing rich semantic information. The sum of this EEG representation and the noise timestep representation (i.e., timestep $t$ in the noise-adding process) encoded by the time embedding layer serves as the conditional input information for the diffusion model. This diffusion model design can utilize the semantically rich external EEG representation for the generative task, leading to a significant improvement in generative performance. Consequently, this enhances the model’s representation learning capability.
@@ III-B The EEG Encoder
The core motivation behind the above EEG encoder design arises from the inherently under-determined nature of the denoising task. The denoiser tends to exploit any relevant knowledge or informative cues to infer and recover the original data distribution as accurately as possible. This incentivizes the encoder to distill salient commonalities between the source and target EEG signals into a compact representation. We constrain the encoded EEG representation to a low-dimensional latent space and guide the denoiser through feature modulation. This strategy encourages the encoder to capture high-level semantic information in the EEG signals, while leaving the reconstruction of local details to the denoiser. Rather than reconstruction from scratch as most conventional denoisers do [hinton2006reducing, higgins2017beta], our strategy enables the encoder to learn a representation that supports latent space denoising. This alleviates the burden of compressing all signal information into the representation, allowing it to prioritize the most distinctive and informative features of the EEG. Such a representation is particularly valuable for adapting to a wide range of downstream EEG tasks.
@@
Our EEG encoder is based on the Vision Transformer (ViT) architecture [dosovitskiy2020image] that operates on the sequences of tokens. We segment each EEG channel into tokens and then flatten the tokens to form consistent biosignal sequences. We retain many of ViT’s best practices, enabling the EEG encoder to be a Transformer architecture that can process input EEG signals with various formats (different sampling rates, number of channels, and time lengths). The input of the encoder is the source EEG signal $x^{\prime}$, which consists of the original samples and the EEG channel augmentations. EEG augmentations are a set of transformations that preserve the semantic information in the EEG channels without altering the interpretation of the EEG data. Here, we employ two augmentations (zero-masking and amplitude scaling) which were shown in prior work [mohsenvand2020contrastive] to be the most effective for extracting useful features.
@@
Specifically, we use a $\omega$-length window without overlap to segment each EEG channel sample into patches ($x^{\prime}=\{x^{\prime}_{j,k}\in\mathbb{R}^{\omega}|j=1,2,\cdots,C,k=1,2,\cdots,{\lfloor{\frac{t^{s}}{\omega}}\rfloor}\}$). The total number of patches for $x^{\prime}$ is ${\left|x^{\prime}\right|}=C{\lfloor{\frac{t^{s}}{\omega}}\rfloor}$. Then a temporal encoder (i.e., a temporal convolution block) is used to encode each EEG patch into a patch embedding $e^{\prime}=\{e^{\prime}_{j,k}\in\mathbb{R}^{d}|j=1,2,\cdots,C,k=1,2,\cdots,{\lfloor{\frac{t^{s}}{\omega}}\rfloor}\}$, where $d$ is the dimension of the embeddings. The learnable positional embeddings are added to all input patch embeddings. Finally, the sequence of embeddings is fed directly into the Transformer network. We use average pooling on the output embeddings to obtain the EEG representation, which is then combined with the noise timestep representation to guide the denoising of the diffusion model. In addition, we perform downstream EEG tasks by using task-specific prediction heads on the EEG representation. As shown in Fig. 2, we train a linear predictor as the prediction head based on the pre-trained encoder in fine-tuning and then analyze the model’s performance on the downstream tasks, in which we employ task-specific loss functions for different downstream objectives. The implementation details can be found in the subsection IV-B.
III-C EEG Data Generation by Diffusion Model